OpenStudy (anonymous):

simplify the expression "-5x^0"

OpenStudy (anonymous):

for every real number \(a\neq 0\)\[a^0=1\]

OpenStudy (anonymous):

so if \(x\neq 0\)\[-5x^0=-5*1=-5\]

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

welcome amigo :)